What is the force of repulsion between two small charges 75.0 cm apart if each has a charge of 1.65 x 10 -12 C? 



1
Expert's answer
2022-05-07T14:40:27-0400
"F=k\frac{q_1q_2}{r^2}\\
=9*10^9*\frac{1.65*10^{-12}*1.65*10^{-12}}{(0.75)^2}=4.36*10^{-14}\rm \: N"